Given square ABCD with AB= 2x+1 and CD= 5x-19 what is the perimeter of the square?: Given square ABCD with AB= 2x+1 and CD= 5x-19 what is the perimeter of the square?
Answer by stanbon(19018)

Given square ABCD with AB= 2x+1 and CD= 5x-19 what is the perimeter of the square?
--------------
All the sides are equal so:
2x+1 = 5x-19
3x = 20
x = 20/3
---------------
One of the sides is 2x+1 = 2(20/3)+1 = 43/3
----------------------
The perimeter = 4(43/3) = 172/3 = 57 1/3
========================
Cheers,
Stan H.